# The Spinning Blimp: Design and Control of a Novel Minimalist Aerial Vehicle Leveraging Rotational Dynamics and Locomotion

### v1

@@ Abstract
This paper presents the Spinning Blimp, a novel lighter-than-air (LTA) aerial vehicle designed for low-energy stable flight. Utilizing an oblate spheroid helium balloon for buoyancy, the vehicle achieves minimal energy consumption while maintaining prolonged airborne states. The unique and low-cost design employs a passively arranged wing coupled with a propeller to induce a spinning behavior, providing inherent pendulum-like stabilization. We propose a control strategy that takes advantage of the continuous revolving nature of the spinning blimp to control translational motion. The cost-effectiveness of the vehicle makes it highly suitable for a variety of applications, such as patrolling, localization, air and turbulence monitoring, and domestic surveillance. Experimental evaluations affirm the design’s efficacy and underscore its potential as a versatile and economically viable solution for aerial applications.
@@ II Vehicle Design
The frame is made from Elmer’s foam core that weighs 224 g per 30$\times$20 sq inch sheet [22]. Coming from a flat sheet, it is laser-cut and folded to form a cage. The frame is attached to the balloon using two elastic strings. Passing through the frame, a square balsa wood stick forms a rigid axle supporting the wings and motors. We designed and evaluated two vehicles with axles of 0.7 m and 1.3 m. The axle length creates effects that have been deeply studied in Blade Element Momentum Theory (BEMT) for a propeller [23, 24].
@@ III Model
The two motors rotation axis being parallel to the $xy$-plane of the body frame. Each motor $i=1,2$ generates a force $f_{i}\geq 0$, located at $\boldsymbol{p}_{1}={[{\ell_{m}},0,0]}$ and $\boldsymbol{p}_{2}={[-{\ell_{m}},0,0]}$ in $\mathcal{B}$. The spinning motion around the $z$-axis of the body frame drags the air using the wings as a propeller, generating a force with magnitude $k_{l}\omega_{z}^{2}$, where $k_{l}$ is a coefficient that abstracts wing parameters such as weight, surface area, and angle of attack.
@@ III-A General dynamics
$\boldsymbol{I}\boldsymbol{\dot{\omega}}+\omega\times\boldsymbol{I}\omega=\boldsymbol{\tau}_{m}+\boldsymbol{\tau}_{bg}+\boldsymbol{\tau}_{d}.$ (2)
$\boldsymbol{\tau}_{m}=\sum_{i=1}^{2}\boldsymbol{p}_{i}\times\boldsymbol{A}\boldsymbol{u}=\ell_{m}\boldsymbol{B}\boldsymbol{u},\text{ where }\boldsymbol{B}=\begin{bmatrix}0&0\\ 0&0\\ 1&1\end{bmatrix}.$
@@ IV Position Control
Figure 5: Control Block Diagram.
The goal of our position control is to move from a position $\boldsymbol{x}=[x,y,z]$ to a desired position $\boldsymbol{x}^{d}$. We control the angular velocity $\omega_{z}$ to maintain a desired height $z^{d}$ while moving towards the goal on the $xy$-plane.
@@ IV-A Height control
The constant $k_{lift}$ was experimentally determined using the bisection method. Then, the height control can be achieved using
@@ IV-A1 Stability in $\omega_{z}$
[…]
@@ IV-C Stability analysis
We use Lyapunov stability to analyze the dynamical system. The Lyapunov conditions are as follows: $V(0)=0,\quad V(x)>0\quad\forall x\neq 0,\quad\dot{V}(x)<0\quad\forall x\neq 0$. To prove stability, we examine stability along the z-axis separately from the x- and y-axes due to the chosen control strategy.
@@ IV-C2 Stability on the $xy$-plane
Without loss of generality, assuming the desired location is at the origin, the direction to the goal is $\beta=\psi+\pi$.
Since $\dot{\boldsymbol{x}}_{xy}=-\dot{\boldsymbol{e}}_{xy}$ and $\ddot{\boldsymbol{x}}_{xy}=-\ddot{\boldsymbol{e}}_{xy}$, the dynamics of the error is
@@
$\dot{V}(\boldsymbol{e}_{xy},\dot{\boldsymbol{e}}_{xy})=\dot{\boldsymbol{e}}_{xy}^{\top}\Big{(}\boldsymbol{e}_{xy}-\frac{2\tau}{m}g(\beta)\begin{bmatrix}\sin\psi\\ -\cos\psi\\ \end{bmatrix}\Big{)}-\frac{1}{m}\begin{bmatrix}d_{x}\,\dot{e}_{x}^{2}\,|\dot{e}_{x}|\\ d_{y}\,\dot{e}_{y}^{2}\,|\dot{e}_{y}|\\ \end{bmatrix}^{\top}.$
@@ V-A Experiment 1: Hover Endurance Test
One of the most relevant design factors of the vehicle is the wing diameter ($2\ell_{w}$). So, we will evaluate short-wings with 0.7m and long-wings with 1.3m. Additionally, we also compare with other LTA vehicles such as the bi-copter [25], and the S-Blimp [20]. Our results show that the endurance for each vehicle is: 30mins for Bi-copter, 75 mins for S-Blimp, 65 mins for Spinning Blimp - short-wing, and the robot with the most endurance is 78 min for the Spinning Blimp with long-wings.

### v2

@@ Abstract
This paper presents the Spinning Blimp, a novel lighter-than-air (LTA) aerial vehicle designed for low-energy stable flight. Using an oblate spheroid helium balloon for buoyancy, the vehicle achieves minimal energy consumption while maintaining prolonged airborne states. The unique and low-cost design employs a passively arranged wing coupled with a propeller to induce a spinning behavior, providing inherent pendulum-like stabilization. We propose a control strategy that takes advantage of the continuous revolving nature of the spinning blimp to control translational motion. The cost-effectiveness of the vehicle makes it highly suitable for a variety of applications, such as patrolling, localization, air and turbulence monitoring, and domestic surveillance. Experimental evaluations affirm the design’s efficacy and underscore its potential as a versatile and economically viable solution for aerial applications.
@@ II Vehicle Design
The frame is made from Elmer’s foam core that weighs 224 g per 30$\times$20 sq inch sheet. Coming from a flat sheet, it is laser-cut and folded to form a cage. The frame is attached to the balloon using two elastic strings. Passing through the frame, a square balsa wood stick forms a rigid axle supporting the wings and motors. We designed and evaluated two vehicles with axles of 0.7 m and 1.3 m. The axle length creates effects that have been deeply studied in Blade Element Momentum Theory (BEMT) for a propeller [22, 23].
@@ III Model
The motors’ rotation axes are parallel to the $xy$-plane of the body frame. Each motor $i=1,2$ generates a force $f_{i}\geq 0$, located at $\boldsymbol{p}_{1}={[{\ell_{m}},0,0]}$ and $\boldsymbol{p}_{2}={[-{\ell_{m}},0,0]}$ in $\mathcal{B}$. The spinning motion around the $z$-axis of the body frame drags the air using the wings as a propeller, generating a force with magnitude $k_{l}\omega_{z}^{2}$, where $k_{l}$ is a coefficient that abstracts wing parameters such as weight, surface area, and angle of attack.
@@ III-A General dynamics
$\boldsymbol{I}\boldsymbol{\dot{\omega}}+\omega\times\boldsymbol{I}\omega=\boldsymbol{\tau}_{m}+\boldsymbol{\tau}_{bg}+\boldsymbol{\tau}_{d},\text{ where}$ (2)
$\boldsymbol{\tau}_{m}=\sum_{i=1}^{2}\boldsymbol{p}_{i}\times\boldsymbol{A}\boldsymbol{u}=\ell_{m}\boldsymbol{B}\boldsymbol{u},\text{ and }\boldsymbol{B}=\begin{bmatrix}0&0\\ 0&0\\ 1&1\end{bmatrix}.$
@@ IV Position Control
Figure 5: Control Block Diagram for position control.
The goal of our position control is to move from a position $\boldsymbol{x}=[x,y,z]$ to a desired position $\boldsymbol{x}^{d}$. We control the angular velocity $\omega_{z}$ to maintain a desired height $z^{d}$ while moving towards the goal on the $xy$-plane (see Fig. 5).
@@ IV-A Height control
The constant $k_{lift}$ is determined experimentally by finding the angular velocity that makes the robot maintain a constant altitude. A systematic way to find the constant is using the bisection method [24]. Then, the height control can be achieved using
@@ IV-A1 Stability in $\omega_{z}$
This first-order differential equation has an exponential solution with $\omega\rightarrow\omega_{z}^{*}$ when time goes to infinity. This provides the necessary condition for the assumption that $\ddot{\psi}=0$ and $\dot{\psi}=\omega_{z}^{*}$ and stability in z-axis.
@@ IV-C Stability analysis
We use Lyapunov stability to analyze the dynamical system. The Lyapunov conditions are as follows: $V(0)=0,\quad V(x)>0\quad\forall x\neq 0,\quad\dot{V}(x)<0\quad\forall x\neq 0$. To prove stability, we analyze stability along the $z$-axis independently of the $x$- and $y$-axes due to the chosen control strategy.
@@ IV-C2 Stability on the $xy$-plane
Without loss of generality, assuming the desired location is at the origin, the direction to the goal is $\beta=\psi+\pi$. Since $\dot{\boldsymbol{x}}_{xy}=-\dot{\boldsymbol{e}}_{xy}$ and $\ddot{\boldsymbol{x}}_{xy}=-\ddot{\boldsymbol{e}}_{xy}$, the dynamics of the error is
@@
$\dot{V}(\boldsymbol{e}_{xy},\dot{\boldsymbol{e}}_{xy})=\dot{\boldsymbol{e}}_{xy}^{\top}\Big{(}\boldsymbol{e}_{xy}-\frac{2\tau}{m}g(\beta)\begin{bmatrix}\sin\psi\\ -\cos\psi\\ \end{bmatrix}\Big{)}-\frac{1}{m}\begin{bmatrix}d_{x}\,\dot{e}_{x}^{2}\,|\dot{e}_{x}|\\ d_{y}\,\dot{e}_{y}^{2}\,|\dot{e}_{y}|\\ \end{bmatrix}^{\top}$
@@ V-A Experiment 1: Hover Endurance Test
One of the most relevant design factors of the vehicle is the wing diameter ($2\ell_{w}$). So, we will evaluate short-wings with 0.7m and long-wings with 1.3m. Additionally, we also compare with other LTA vehicles such as the bi-copter [25, 26], and the S-Blimp [20]. Our results show that the endurance for each vehicle is: 30mins for Bi-copter, 75 mins for S-Blimp, 65 mins for Spinning Blimp - short-wing, and the robot with the most endurance is 78 min for the Spinning Blimp with long-wings.